OpenStudy (anonymous):

Solve the equation \[\sqrt{2x-3}\] +5 =6

OpenStudy (primeralph):

draw it.

OpenStudy (anonymous):

\[\sqrt{2x-3}\] +5 =6 like this

hartnn (hartnn):

you can subtract 5 fom both sides, first

OpenStudy (anonymous):

There's no need to draw it though

OpenStudy (anonymous):

but its not underneath its next to it

hartnn (hartnn):

then squaring both sides will make it much simpler then

OpenStudy (anonymous):

im so confused

hartnn (hartnn):

isolate the square root sign by subtracting 5 from both sides...can u?

OpenStudy (anonymous):

yea i did that and got 1

hartnn (hartnn):

correct...now square both sides..

OpenStudy (anonymous):

ok and i got 4x²-20x+25

hartnn (hartnn):

actually, squaring and square root gets cancelled \((\sqrt{A})^2=A\) so, \((\sqrt{2x-3})^2=2x-3\) only got this ?

OpenStudy (anonymous):

I will show you how it's done just now that is if you will like how I will do it but even if you don't like my way I will do it my way anyway : \[\sqrt{2x-3}+5=6 \rightarrow \sqrt{2x-3}=6-5 \rightarrow \sqrt{2x-3}=1 \] And now you can square both sides of the equation as follows \[\left( \sqrt{2x-3} \right)^2 = 1^2 \] \[2x-3=1\] \[2x=4 \rightarrow x=2\] but you need to go back and check if the solution makes sense or not hey by substituting your result into the original equation \[\sqrt{2(2)-3} +5=6\] \[\sqrt{4-3}+5=6 \rightarrow \sqrt{1}+5=6 \rightarrow 1+5=6 \] and you can see that hey the left hand side is equal to the right hand side which means that the solution makes sense \[6=6 \rightarrow RHS=LHS\] as expected

OpenStudy (anonymous):

kinda?

hartnn (hartnn):

so its just 2x-3 =1 aon't this easy to solve ?

hartnn (hartnn):

*ain't

OpenStudy (anonymous):

yea i its x=2

hartnn (hartnn):

correct :)

OpenStudy (anonymous):

Does the answer make sense to you though ? Indiandabest ???

OpenStudy (anonymous):

i mean kinda i kinda know what to do from now on

OpenStudy (anonymous):

Dont I deserve a medal for this at least :)

OpenStudy (anonymous):

lol yea

OpenStudy (anonymous):

Hahaha thanks